# The accreted stellar haloes of Milky Way-mass galaxies as a probe of the nature of the dark matter

### v1

@@ 4.3 Kinematics
Figure 3: Radial velocity of accreted stellar particles, as a function of their radial distance to the centre of their host galaxies. The colour of each particle encodes the local ex-situ stellar density. This is shown for three $z=0$ haloes (top to bottom) across their matched CDM, WDM and SIDM versions (left to right). The vertical ordering reflects the ease with which differences can be identified across DM models, with their IDs (G1, G2 and G3) shown on the right hand side. The approximate time at which each galaxy experienced its last major merger is also indicated on the right-hand axis, which is similar across DM models (see Table 1).
@@
To investigate how the velocity distribution of the ex-situ stellar halo changes across DM models, we first compare three representative galaxies from our sample. They were chosen based on the fact that no disruptions of massive satellites have occurred in the recent past, providing a view into how a relaxed stellar halo looks like. We show in Fig. 3 how the radial velocity of ex-situ stars, $v_{r}$, vary as a function of their distance to the centre, $r$. From hereon, we refer to these examples as G1, G2 and G3, which is intended to represent the relative ease with which differences in the kinematics of accreted stars can be visually identified across DM models. The $v_{r}$ vs $r$ distributions for the rest of the MW-mass haloes in our sample are shown in Fig. 8.
@@
Several ‘chevrons’ – i.e. overdensities in $r,v_{r}$ space – are present in G1 and G2, whereas the distribution of G3 is largely smooth and featureless. Chevrons reflect apocentre pileups of stars that have been stripped from their progenitor galaxy, so their number and location encode information about the stripping history of the progenitor satellite that gave rise to them (e.g. Dong-Páez et al., 2022). The subsequent evolution is governed by interactions with the local environment (e.g. Davies et al., 2023a, b) and Liouville’s theorem, whereby chevrons progressively wind up in position and velocity space until they can no longer be distinguished from one another nor the background.
@@
Focusing on the first two examples, it is clear that the number of chevrons varies according to the assumed dark matter model, with WDM having the fewest and SIDM having the most. The thickness of each individual chevron also changes with the assumed dark matter model, with the thinnest chevrons being present in SIDM. In line with the spatial distribution of the accreted stellar halo, the SIDM chevrons are typically found at larger radii than in the other two models. Since this entails longer dynamical timescales, the time for the chevrons to fold and mix with the background is expected to be longer in SIDM than in WDM and CDM.
@@
To put these examples in perspective with respect to the rest of the sample of stellar haloes used in this study, we highlight with green circles their ex-situ masses in Fig. 1. Their masses bracket that of the overall sample, and the ordering based on the visual comparison of the differences in the number and prominence of chevrons reflects their mass ranking. Since the mass of stellar haloes typically correlate to their formation time, we proceed to identify when the stellar haloes of G1, G2 and G3 formed.
@@
Table 1: Summary properties of the last major mergers experienced by the three examples shown in Fig. 3, identified independently across the DM variants considered in this work. The infall time, $t_{\mathrm{infall}}$, corresponds to when the merged object first became a satellite of the main progenitor. We measure the merger mass ratio in the output immediately prior to this time, before significant tidal stripping occurs. The merger is complete at $t_{\mathrm{merge}}$, which we identify as the time when the object no longer had a resolved bound component. The time elapsed between infall and accretion is shown in the rightmost column.
Our operational definition for the formation time of the accreted stellar halo is based on when the last major merger occurred. To identify when this happened, we follow the merger tree of every MW-mass halo in our sample, and identify all objects that directly merged onto its main progenitor branch. We then compute the mass ratio between the merged object and the main progenitor, $f_{\mathrm{merge}}$, at the time before the stellar halo progenitor became a satellite of the host for the first time. This avoids underestimating the merger mass ratio due to subsequent mass loss and growth of the host. In practice, we use the FoF group membership to identify when the stellar halo progenitor and the host main progenitor are in the same FoF group, which we use as our definition of accretion. Finally, we select all luminous objects with $f_{\mathrm{merge}}\geq 0.1$, and inspect their evolution to check whether it was indeed a major merger. If no candidates are found, we progressively lower the threshold until one is identified. This was only needed for example G3, which formed early and hence had a less trivial analysis.
The resulting formation times of the stellar haloes of G1, G2 and G3 are $t_{\mathrm{form}}\equiv t_{\mathrm{infall}}\sim 5.78,4.71,1.90~{}\mathrm{Gyr}$, respectively. There is little to no variation in when these occur across DM models, aside from G3, which we attribute to the aforementioned difficulty associated to its early forming nature. By identifying the formation time, we see that the magnitude of the differences in the abundance and prominence of chevrons across DM models reflects the age of the accreted stellar halo, with the largest changes present in later forming stellar haloes. Even if differences existed shortly after the last major merger occurred, the subsequent evolution of chevrons makes any initially existing ones mix beyond our visual identification capabilities.
@@
Figure 4: Top panel: velocity anisotropy of the ex-situ stellar component around the MW-mass haloes in our sample, across CDM (black), WDM (red) and SIDM (blue) models. The shaded regions indicate the 16th - 84th percentiles and the solid lines the median values. Both have been smoothed using a linear Savitzky–Golay filter over three consecutive bins. Bottom panel: same as above, but showing the median difference in the velocity anisotropy across matched counterparts, and the associated 16th - 84th percentiles.
@@
These differences propagate into the merging time of the progenitors, which we define to be the difference in time between their accretion and disruption [^3] Defined to occur when SUBFIND is unable to identify them as self-bound structure during five consecutive outputs.. As shown in Table 1, they vary across models in a systematic manner: SIDM subhaloes disrupt the fastest, and the WDM ones take the longest to do so. The change in SIDM is likely due to the presence of flat density cores and subhalo evaporation, which exacerbates mass loss and accelerates their disruption. In WDM, their apocentres are larger than in CDM after the first pericentric passage. This could be due to their aforementioned (small) differences in their location in $r,v_{r}$ space relative to CDM, which are amplified after undergoing pericentre. Consequently, they remain further out and their disruption takes longer. Whether this is a systematic difference always present in accreted WDM progenitors of stellar haloes, or a difference only present in the three examples discussed here, is relegated to future work.
@@
The primary progenitors of the stellar haloes of G1, G2 and G3 have high mass ratios relative to the main progenitor of the MW-mass hosts (see Table 1). Their post-accretion evolution is therefore driven by the interplay between dynamical friction and loss of bound mass. As their masses at accretion time do not differ substantially across DM models, changing how long mergers last for could affect the integrated effect they experience from dynamical friction. Hence, the systematically different merging timescales suggest that WDM and SIDM progenitors might have experienced more or less net dynamical friction than the corresponding CDM counterparts. This could explain why the spatial distributions end up being slightly denser in WDM and substantially less dense in SIDM, respectively. However, as noted before, the ex-situ density changes occur on the same scale as the DM ones in SIDM. This highlights the fact that the end result could be a combination of changing dynamical friction, subhalo stripping & evaporation, and the gravitational potential of the host. Identifying which of these factors dominates is not trivial to do in the cosmological simulations we use in this work.
@@
Galaxies that experience major mergers, as is the case during the formation of the stellar haloes of G1, G2 and G3, typically undergo orbit radialisation (e.g. Amorisco, 2017; Vasiliev et al., 2022). This process removes angular momentum from the orbits of merging galaxies, making them more radial than when they were first accreted. Given the differences in how long it took for the progenitors in G1, G2 and G3 to merge across DM models, the anisotropy of accreted stars could, in principle, also change. To investigate whether this is the case, we measure the velocity anisotropy of all ex-situ stars, defined as:
@@
across all haloes in our sample. We use a coordinate system where the angular momentum of all stellar particles within 5 kpc of the centre points along the z-axis. We use a similar binning scheme as the density profiles, i.e., 20 spherical shells logarithmically spaced between $0.005R_{200}$ and $R_{200}$, and then take the median values across our whole sample.
@@
The fact that such a difference exists is striking, and highlights the potential usefulness of the Gaia-Sausage-Enceladus (GE/S) debris as a DM constraint. In other words, given current constraints on its orbital anisotropy (e.g. $\beta\sim 0.8$; Belokurov et al., 2018), one can identify how large of an SIDM cross-section can be accommodated before radialisation becomes ineffective and no GE/S-like remnants are produced. This would provide hints on what the SIDM cross-section on scales between MW-mass and dwarf-mass scales is, corresponding to the transition region for velocity-dependent cross-sections.
Figure 5: Top panel: distribution of the peak maximum circular velocity of the galaxies that directly merge with the MW-mass main progenitors, which we take as the building blocks of their stellar halo. This is shown across different DM models, as indicated by the top right legend. The solid lines indicate the median values across our population, with the shaded regions showing the 16th and 84th percentiles of the distribution. Bottom panel: same as above, but for the median difference across matched counterparts. See Fig. 7 for the corresponding peak stellar mass distributions.
@@
4.4 The building blocks
Figure 6: Cumulative distribution function of the pericentric distances of the stellar halo progenitors before they disrupted. They are classified into two $V_{\mathrm{peak}}$ bins, according to whether they have values larger (right column) or lower (left column) than $50~{}\mathrm{km}\,\mathrm{s}^{-1}$. Top panel: median distribution, with shaded regions bracketing the 16th and 84th percentiles. Middle panel: same as above, but expressed relative to the value of their matched CDM counterpart. The dashed line indicates the scale at which at least one CDM counterpart reaches a value of zero, and hence where the median ratios start becoming dominated by fewer objects. Bottom panel: same as the two panels above, but normalised to the total number of objects within $R_{200}$. All these distributions have been smoothed using a linear Savitzky–Golay filter over three consecutive bins.

### v2

@@ 4.3 Kinematics
Figure 3: Radial velocity of accreted stellar particles, as a function of their radial distance to the centre of their host galaxies. The colour of each particle encodes the local ex-situ stellar density. This is shown for three $z=0$ haloes (top to bottom) across their matched CDM, WDM and SIDM versions (left to right). The vertical ordering reflects the ease with which differences can be identified across DM models, with their IDs (G1, G2 and G3) shown on the right hand side. The infall time of the satellite galaxy that was involved in the last major merger for each of the galaxies in this sample is indicated on the right-hand side of the corresponding row. The values are similar across DM models (see Table 1). The plots below each distribution show the density of ex-situ stars with $|v_{r}|\leq 50\,\mathrm{km}\,\mathrm{s}^{-1}$ measured using 1 kpc radial bins, relative to the smoothed density found using a 6 kpc running mean. The solid black line and the shaded region correspond to the median and 2nd to 98th percentiles, respectively. The vertical red lines indicate peaks whose 2nd percentile values satisfy $\rho/\langle\rho\rangle>1$, which correspond to the location of the most prominent chevrons.
We now explore whether or not the velocity distribution of the ex-situ stellar halo changes across DM models. We first compare how the radial velocities of ex-situ stars, $v_{r}$, vary as a function of their distance to the centre, $r$, for three illustrative galaxies from our sample (§4.3.1). Following the discussion of the three illustrative galaxies, we shift our focus to the whole sample of MW-mass galaxies to investigate whether or not the velocity anisotropy of the ex-situ stars changes with the DM model in a systematic manner (§4.3.2).
4.3.1 The phase-space and build up of the stellar halo around three illustrative galaxies
@@
The three galaxies we discuss are chosen based on the fact that their $v_{r}$ vs $r$ distributions are not significantly asymmetrical, indicating that no significant disruptions of satellites have occurred in the recent past and thus provide a view of what a relaxed stellar halo looks like. From hereon, we refer to these examples as G1, G2 and G3, which is intended to represent the relative ease with which differences in the kinematics of accreted stars can be visually identified across DM models.
We show in Fig. 3 the $v_{r}$ vs $r$ distribution of the ex-situ stars for the three galaxies. The $v_{r}$ vs $r$ distributions for the other five MW-mass haloes in our sample are shown in Fig. 8, which show a greater variety in their distributions than the three examples we discuss in this section. Some galaxies, like G4, have significant asymmetries as a result of the ongoing disruption of satellites. Other galaxies, such as G5 and G6, have a mostly featureless and smooth distribution, aside from one or two overdensities in $r,v_{r}$ space (‘chevrons‘). The remaining two galaxies (G7 and G8) have prominent chevrons, which tend to follow the trends we identify using G1, G2 and G3 as illustrative examples.
@@
Several chevrons are present in G1 and G2, whereas the distribution of G3 is largely smooth and featureless. Chevrons reflect apocentre pileups of stars that have been stripped from their progenitor galaxy, so their number and location encode information about the stripping history of the progenitor satellite that gave rise to them (e.g. Dong-Páez et al., 2022). The subsequent evolution is governed by interactions with the local environment (e.g. Davies et al., 2023a, b) and Liouville’s theorem, whereby chevrons progressively wind up in position and velocity space until they can no longer be distinguished from one another nor the background.
@@
Focusing on G1 and G2, it is clear that the number of chevrons varies according to the assumed dark matter model, with WDM having the fewest and SIDM having the most. The thickness of each individual chevron also changes with the assumed dark matter model, with the thinnest chevrons being present in SIDM. In line with the spatial distribution of the accreted stellar halo, the SIDM chevrons are typically found at larger radii than in the other two models. Since this entails longer dynamical timescales, the time for the chevrons to fold and mix with the background is expected to be longer in SIDM than in WDM and CDM.
We also show in Fig. 3 how the local mass density of ex-situ stars with low radial velocities compares to the background ex-situ stellar mass density. Since stars that are part of chevrons resemble Gaussian distributions at low $v_{r}$ values (e.g. Donlon et al., 2024), the local ex-situ density at the pile-up location is higher than neighbouring radial bins. Thus, measuring the contrast between the local and background ex-situ stellar densities for low $v_{r}$ can help identify the number and location of chevrons.
In practice, we measure the local mass density of stars with $|v_{r}|\leq 50\,\mathrm{km}\,\mathrm{s}^{-1}$ using 1 kpc radial bins. We then estimate the background ex-situ density by smoothing the local densities using a running mean with a window length of 6 kpc. The smoothing is done to remove the radial dependence of the ex-situ stellar density. As the density contrast can be prone to noise, we estimate the 2nd, 50th and 98th percentile values at each radial bin after performing 20000 bootstrap resamplings of the low radial velocity stellar particles. After estimating the spread in values of the density ratio as a function of radius, we identify the location of statistically significant peaks [^4] The number of statistically significant peaks is sensitive to the percentile threshold used to identify them, as well as the window length used to estimate the background ex-situ density. Using less stringent percentile value thresholds identifies more peaks that correspond to visually-identifiable chevrons, but also more peaks purely driven by noise fluctuations. Smoothing over smaller scales leads to more prominent peaks, including those driven by noise. Conversely, computing the average over longer spatial scales can smooth out the thinnest peaks caused by chevrons. based on which radial bins have a 2nd percentile value above unity. If more than one consecutive radial bin satisfies this threshold, we consider them to be part of the same peak and take as its position the average of the radial bin centres. This approach results in the red vertical lines shown in the $\rho/\langle\rho\rangle$ panels of Fig. 3.
Beyond the inner 5 kpc of the galaxy, the location and number of statistically significant peaks in the local density contrast correlate well with those of the most prominent chevrons. For G1, the SIDM version contains the largest amount of peaks, which are typically $\approx 30\%$ denser than the background in which they are embedded. Some troughs are also present at a radius immediately adjacent to the edge of a well-defined chevron. In some cases, the density contrast of troughs falls below the unity line, which is caused by the smoothing of the local density to estimate the background density. In contrast, the CDM and WDM versions of G1 do not have as many statistically significant peaks as the SIDM version, and those which are clearly present do not reach density contrasts as high as those found in SIDM. Repeating the analysis on G2 and G3, we see that the number of statistically significant density peaks reduces even further, and that the estimates get noisier due to fewer particles sampling the stellar halo.
@@
To put these examples in perspective with respect to the rest of the sample of stellar haloes used in this study, we highlight with green circles their ex-situ masses in Fig. 1. Their masses bracket that of the overall sample, and the ordering based on the visual comparison of the differences in the number and prominence of chevrons reflects their mass ranking. Since the mass of stellar haloes typically correlate to their formation time, we proceed to estimate when the stellar haloes of G1, G2 and G3 finished accreting most of their mass.
@@
Table 1: Summary properties of the last major mergers experienced by the three examples shown in Fig. 3, identified independently across the DM variants considered in this work. The infall time, $t_{\mathrm{infall}}$, corresponds to when the merged object first became a satellite of the main progenitor. We measure the merger mass ratio in the output immediately prior to this time, before significant tidal stripping occurs. The merger is complete at $t_{\mathrm{merge}}$, which we identify as the time when the object no longer had a resolved bound component. The time elapsed between infall and disruption is shown in the rightmost column.
For this purpose, we identify for each galaxy in our sample when the its last major merger was accreted. To find the infall time of the last major merger, we follow the main branch of our G1, G2 and G3 example galaxies. We subsequently identify all galaxies that directly merged onto the main progenitor branch of either of these three example galaxies. To determine whether the accretion of a given galaxy was associated with a major merger, we compute the mass ratio between the merged galaxy and the main progenitor, $f_{\mathrm{merge}}$, at the time before the stellar halo progenitor became a satellite of the host for the first time [^5] This definition of mass ratio avoids underestimating the merger mass ratio due to post-infall mass loss and growth of the host.. In practice, we use the FoF group membership to identify when the stellar halo progenitor and the host main progenitor are in the same FoF group, which we use as our definition of infall. Finally, we select all luminous objects with $f_{\mathrm{merge}}\geq 0.1$, and inspect their evolution to check whether it was indeed a major merger. If no candidates are found, we progressively lower the threshold until one is identified. This was only needed for example G3, which formed early and hence had a less trivial analysis.
The resulting infall times of the galaxy involved in the last major merger for each of the G1, G2 and G3 galaxies are $t_{\mathrm{infall}}\approx 5.78,4.71,1.90~\mathrm{Gyr}$, respectively. There is little to no variation in when these occur across DM models, aside from G3, which we attribute to the aforementioned difficulty associated to its early forming nature. We see that the magnitude of the differences in the abundance and prominence of chevrons across DM models correlates with the time when the last major accretion took place, with the largest changes present in galaxies that experienced later major mergers. Even if differences existed shortly after the last major merger occurred, the subsequent evolution of chevrons makes any initially existing ones mix beyond our visual identification capabilities.
@@
Figure 4: Top panel: velocity anisotropy of the ex-situ stellar component, averaged across the eight MW-mass haloes in our sample in their CDM (black), WDM (red) and SIDM (blue) versions. The shaded regions indicate the 16th - 84th percentiles and the solid lines the median values. Both have been smoothed using a linear Savitzky–Golay filter over three consecutive bins. Bottom panel: same as above, but showing the median difference in the velocity anisotropy across matched counterparts, and the associated 16th - 84th percentiles.
@@
These differences propagate into the merging time of the progenitors, which we define to be the difference in time between their accretion and disruption [^6] Defined to occur when SUBFIND is unable to identify them as self-bound structure during five consecutive outputs.. As shown for the three example galaxies in Table 1, they vary across models in a systematic manner: SIDM subhaloes disrupt the fastest, and the WDM ones take the longest to do so. The change in SIDM is likely due to the presence of flat density cores and subhalo evaporation, which exacerbates mass loss and accelerates their disruption. In WDM, their apocentres are larger than in CDM after the first pericentric passage. This could be due to their aforementioned (small) differences in their location in $r,v_{r}$ space relative to CDM, which are amplified after undergoing pericentre. Consequently, they remain further out and their disruption takes longer. Whether this is a systematic difference always present in accreted WDM progenitors of stellar haloes, or a difference only present in the three examples discussed here, is relegated to future work.
@@
The primary progenitors of the stellar haloes of G1, G2 and G3 have high mass ratios relative to the main progenitor of the MW-mass hosts (see Table 1). Their post-accretion evolution is therefore driven by the interplay between dynamical friction and loss of bound mass. As their masses at infall time do not differ substantially across DM models, changing how long mergers last for could affect the integrated effect they experience from dynamical friction. Hence, the systematically different merging timescales suggest that WDM and SIDM progenitors might have experienced more or less net dynamical friction than the corresponding CDM counterparts. This could explain why the spatial distributions end up being slightly denser in WDM and substantially less dense in SIDM, respectively. However, as noted before, the ex-situ density changes occur on the same scale as the DM ones in SIDM. This highlights the fact that the end result could be a combination of changing dynamical friction, subhalo stripping & evaporation, and the gravitational potential of the host. Identifying which of these factors dominates is not trivial to do in the cosmological simulations we use in this work.
4.3.2 The velocity anisotropy profile
The most massive progenitors of stellar haloes generally have non-negligible masses compared to their host galaxy at infall time, as we have just illustrated using the examples of G1, G2 and G3. The most massive progenitors are therefore likely to experience a major merger, which is typically characterised by dynamical friction and orbit radialisation (e.g. Amorisco, 2017; Vasiliev et al., 2022). Orbit radialisation, which results from the removal of angular momentum from the orbits of merging galaxies, makes their orbits more radial than when they were first accreted.
@@ 4.3.2 The velocity anisotropy profile
Motivated by the differences in how long it took for the progenitors in G1, G2 and G3 to merge across DM models, the anisotropy of accreted stars could, in principle, also change in response to varying degrees of orbit radialisation. To investigate whether this is the case, we measure the velocity anisotropy of all ex-situ stars, defined as:
@@
across all eight MW-mass haloes in our sample. We use a similar binning scheme as the density profiles, i.e. 20 spherical shells logarithmically spaced between $0.005R_{200}$ and $R_{200}$, and then take the median values across our whole sample.
We measure the velocity anisotropy of the ex-situ stellar component using a galactocentric coordinate system, defined independently for each of the eight galaxies in our sample. The plane of the galaxy is defined as being perpendicular to the mass-weighted angular momentum vector of all stars within 5 kpc of its centre. We visually confirm that this procedure results in a good alignment with the stellar disc, if it is present. Two of the galaxies in our sample do not have clearly identifiable discs and have a more spheroidal stellar distribution.
@@
The fact that such a difference exists is striking and highlights the potential usefulness of the Gaia-Sausage-Enceladus (GE/S) debris as a DM constraint. In other words, given current constraints on its orbital anisotropy (e.g. $\beta\approx 0.8$; Belokurov et al., 2018), one can identify how large of an SIDM cross-section can be accommodated before radialisation becomes ineffective and no GE/S-like remnants are produced. This would provide hints on what the SIDM cross-section on scales between MW-mass and dwarf-mass scales is, corresponding to the transition region for velocity-dependent cross-sections.
Exploring the feasibility of this new potential constrain will require a more dedicated analysis than we have done here, which could take on several forms. One could be a targeted approach, whereby one identifies in a CDM simulation which MW-mass galaxies have a GE/S-like debris and see by how much its velocity anisotropy and the evolution of the GE/S progenitor differs in its SIDM counterparts. A more general approach would be to ask the question of how probable it is to find a highly radially anisotropic GE/S-like debris in different SIDM models, without having to rely on matching to CDM simulations. The benefit of not relying on CDM simulations to select the SIDM counterparts to study is that the population of galaxies and set of assembly histories that give rise to GE/S-like debris may differ between CDM and SIDM.
Figure 5: Top panel: distribution of the peak maximum circular velocity of the galaxies that directly merge with the main progenitors of the eight MW-mass haloes of our sample, which we take as the building blocks of their stellar halo. This is shown across different DM models, as indicated by the top right legend. The solid lines indicate the median values across our population, with the shaded regions showing the 16th and 84th percentiles of the distribution. Bottom panel: same as above, but for the median difference across matched counterparts. See Fig. 7 for the corresponding peak stellar mass distributions.
@@
4.4 The properties of stellar halo progenitors
Figure 6: Cumulative distribution function of the pericentric distances of the stellar halo progenitors before they disrupted, measured for the eight MW-mass haloes in our sample. They are classified into two $V_{\mathrm{peak}}$ bins, according to whether they have values larger (right column) or lower (left column) than $50~\mathrm{km}\,\mathrm{s}^{-1}$. Top panel: median distribution, with shaded regions bracketing the 16th and 84th percentiles. Middle panel: same as above, but expressed relative to the value of their matched CDM counterpart. The dashed line indicates the scale at which at least one CDM counterpart reaches a value of zero, and hence where the median ratios start becoming dominated by fewer objects. Bottom panel: same as the two panels above, but normalised to the total number of objects within $R_{200}$. All these distributions have been smoothed using a linear Savitzky–Golay filter over three consecutive bins.